OpenStudy (anonymous):

STATISTICS! Hypothesis Testing: ‘There is no difference between the products’ effectiveness’. • Using the variable Score A – Score B, conduct a two-tailed Z-test at the 5% significance level to determine if there is really a statistically significant difference between the effectiveness of each mouthwash. Check that the conditions are met to use the z-test. • Calculate the 95% confidence interval for the difference between the scores. • The dental student wrote a report on their findings concluding that “significantly less plaque accumulation will result from using Brand B mouthwas

OpenStudy (anonymous):

See table below: Eight of the volunteers were given mouthwash A firstand then mouthwash B second, for days, The other eight volunteers were given mouthwash B first and then mouthwash A second.

OpenStudy (anonymous):

My question is regarding finding the confidence interval. I am not sure which proportion to use in the formula...

OpenStudy (anonymous):

\[p ^{*} -1.96\sqrt{\frac{p ^{*}(1-p ^{*}) }{ n }} \le p \le p ^{*} +1.96\sqrt{\frac{p ^{*}(1-p ^{*}) }{ n }}\]

OpenStudy (paki):

hey... please post your question in the relevant section, so that it can be guided/answered properly... thanks...

OpenStudy (anonymous):

Sorry! I made a mistake... will do :)

OpenStudy (paki):

thank you...